# Maud Biquard ISAE-Supaero / CNES

### v1

@@ 4.1 Variational and compressive autoencoders
$\mathcal{L}(\theta,\phi,\gamma;x)=\mathbb{E}_{q_{\phi}(z|x)}\left[\frac{1}{2\gamma^{2}}||x-D_{\theta}(z)||^{2}_{2}\right]$
@@
Compressive autoencoders (CAEs) are neural networks used for compression, which present similarities with VAEs. CAEs are trained by optimizing the so-called rate distortion trade-off = $\mathcal{L}=\mbox{ Rate}+\alpha\times\mbox{Distortion}$. The distortion represents an error between the input and output of the CAE, and the rate measures the compression efficiency. CAEs can be expressed as VAEs with specific inference and generative distributions [2]. In particular, the encoder posterior $q_{\phi}(z|x)$ is a uniform distribution, of mean $\bar{z}$ and with a fixed variance in order to simulate the quantization process. Compared to VAE loss, the distortion resembles the data fidelity term, while the rate corresponds to the latent constraint. The trade-off parameter $\alpha$ plays a similar role as $\gamma$ for VAEs, controlling the trade-off between the two loss terms.
Besides, state-of-the-art CAEs use an hyperprior module [2, 3, 11], as described in Fig. 1. This additional autoencoder takes $\bar{z}$ as input and predicts $z$’s mean $\mu^{z}$ and deviation $\sigma^{z}$. The latent prior on $z$ is then defined as the factorized distribution $\prod_{k}\mathcal{N}(\mu^{z}_{k},(\sigma^{z}_{k})^{2})$. Hence, the latent prior is a $z$-adaptive distribution, more powerful and flexible than the traditional VAE prior $\mathcal{N}(0,I)$. Formally, these CAEs can be expressed as VAEs with two latent variables $z$ and $h$ [3].
@@ 4.2 Background on VBLE with CAEs
Figure 1: A compressive autoencoder (CAE) with a hyperprior.
Variational Bayes Latent Estimation (VBLE) [5] is an algorithm used for image restoration, which enables efficient approximate posterior sampling. It belongs to latent optimization methods [6], which estimate the solution of an inverse problem in the latent space of a generative model. In place of a classical generative model, [5] proposed to use CAEs trained on image compression tasks.
Furthermore, the VBLE method estimates the latent inverse problem posterior $p_{Z|Y}(z|y)$ in the latent space of the CAE using variational inference. Specifically, it considers the following factorized parametric family:
$E_{\bar{z},a}$ =$=\left\{q_{\bar{z},a}(z)\big{|}\bar{z},a\in\mathbb{R}^{C\times M\times N},a>0\right\}$ (13)
[…]
@@
$\mathcal{L}(\bar{z},a;y)=\mathbb{E}_{q_{\bar{z},a}(z)}\big{[}\log p_{Y|Z}(y|z)+\log p_{\theta}(z)$
$-\log q_{\bar{z},a}(z)\big{]}$ (14)
where $p_{\theta}(z)$ is the latent prior, that is the hyperprior $\mathcal{N}(\mu^{z},\sigma^{z})$ in the case of CAEs, with the notation of Fig. 1. Using the reparameterization trick $z=\bar{z}+au$ with $u\sim\mathcal{U}(-\frac{1}{2},\frac{1}{2})$ [38], eq. 14 can be optimized with a Stochastic Gradient Variational Bayes (SGVB) algorithm [37].
@@ 4.3 Limitation of latent posterior estimation
VBLE with CAEs [5] enables efficient posterior sampling. Therefore, it seems of great interest for satellite image restoration as employing heavier methods would hardly be scalable. However, having reliable uncertainties is necessary if these are to be used in practical applications.
In our experiments, VBLE tends to underestimate uncertainties. To illustrate this, the first line of Tab. 1 provides confidence interval prediction probabilities (ICP) at level 95% using VBLE with a simple VAE on CelebA. Precisely, for each pixel of an image, we compute a confidence interval of a given level using 100 posterior samples. Then, ICP denotes the proportion of the ground truth pixels lying in the predicted intervals. Table 1 first line shows too low ICPs indicating that uncertainty predictions could be improved.
$\sigma=2.55/255$ $\sigma=7.65/255$
CelebA 0.84 0.93
ManifoldCelebA 0.92 0.96
Table 1: Interval coverage probability of level 95% for two deblurring problems with Gaussian noise level $\sigma$. Averaged on 20 test images.
Yet, in latent optimization methods, there are three types of errors [6]: the optimization error, the inverse problem error (caused by the inverse problem’s ill-posedness), and the representation error (when the solution is not in the generator range). While the first two errors can reasonably be modelled by the latent posterior in VBLE, the representation error lies inherently in the image space, thus is hardly taken into account in the latent space. Therefore, we hypothesize that VBLE underestimates the uncertainties mainly because of this source of error. In Tab. 1, the second line corresponds to ICP results with a dataset that we called ManifoldCelebA. This dataset consists of CelebA images which have been encoded and decoded with the VAE, so that the representation error is null. ICPs for ManifoldCelebA are much closer to 0.95, validating our hypothesis.
4.4 Proposed joint latent and image posterior model
In the following, we propose a new method for image restoration, derived from VBLE, in which we model the inverse problem posterior jointly in the latent and image space in order to better model the representation error. For the sake of simplicity, we only consider one latent variable $z$, but the equation can easily be adapted for two latent variable CAEs by considering a concatenated variable $\psi=(z,h)$ and $p(\psi)=p(z|h)p(h)$.
Figure 2: Generative model graph of our variational inference algorithm.
We aim to estimate the joint latent and image inverse problem posterior $p_{X,Z|Y}(x,z|y)$ using variational inference. Note that for VBLE, the targeted distribution was the latent posterior $p_{Z|Y}(z|y)$. For that purpose, we consider the following parametric family, corresponding to the representation of Fig. 2:
$E_{\bar{z},a,b}$ =$=\{q_{b}(x|z)q_{\bar{z},a}(z)|\bar{z},a\in\mathbb{R}^{C\times M\times N},$
@@ 4.4 Proposed joint latent and image posterior model
with $C^{\prime},M^{\prime}$ and $N^{\prime}$ being the channel and spatial dimensions of the image space, $\sigma_{\theta}(z)$ being the deviation of the generative model $p_{\theta}(x|z)=\mathcal{N}(D_{\theta}(z),\Sigma_{\theta}(z):=\operatorname{diag}(\sigma_{\theta}(z)^{2})$. This parametric distribution is composed of a latent term $q_{\bar{z},a}(z)$ and of an image term $q_{b}(x|z)$ which was designed to take into account the representation error. As in VBLE, the latent term is chosen to be uniform as it matches the CAE encoder distribution. The choice for $q_{b}(x|z)$ is motivated by the fact that $p_{\theta}(x|z)$ represents the likelihood of generating $x$ given $z$. Hence, even if the representation error for an image $x^{*}$ is nonzero (that is $D_{\theta}(z)\neq x^{*}$), $x^{*}$ should nevertheless lie in a zone where $p_{\theta}(x|z)$ is high. Then, parameter $b$ allows to better adapt the posterior model to the given inverse problem.
@@
$+\log p_{\theta}(x|z)+\log p_{\theta}(z)-\log q_{b}(x|z)-\log q_{\bar{z},a}(z)\big{]}$ (15)
with $p_{\theta}(z)$ being the latent prior of the generative model, that is the hyperprior in the CAE case. Then, the $(x|z)$ terms represent a KL divergence between two Gaussian distributions and can be easily simplified, as well as $\log q_{\bar{z},a}(z)$ which is independent of $z$. Thus:
@@
$+\log p_{\theta}(z)\big{]}+\sum_{k}\log a_{k}+\sum_{i}\log b_{i}-\frac{1}{2}b_{i}^{2}.$ (16)
@@
eq. 16 can be optimized using gradient descent with a Monte Carlo approximation of the ELBO gradient as in [38, 5].
4.5 Learning the generator variance
The algorithm presented in the last section requires the following generative model: $p_{\theta}(x|z)=\mathcal{N}(D_{\theta}(z),diag(\sigma_{\theta}(z)^{2})$, thus the generator should estimate a pixelwise variance $\sigma_{\theta}(z)$. However, in classical VAE and CAE models, the decoder distribution is a Gaussian distribution $p_{\theta}(x|z)=\mathcal{N}(D_{\theta}(z),\gamma^{2}I)$ with a fixed perpixel standard deviation $\gamma$ responsible for the trade-off between the data fidelity and the latent constraint terms in the loss. This distribution is not appropriate in our case as the reconstruction error should be larger on high frequency areas than on large flat ones.
Some approaches learn more representative generator distributions [58, 19, 17]. Following [58], we add a second decoder predicting $\sigma_{\theta}(z)$, which has the same structure as the original decoder $D_{\theta}$. We adopt a two stage training strategy as is typically done for training stability: first, the CAE is classically trained with a given $\gamma$ trade-off parameter. Then, the second decoder is trained by minimizing the negative log-likelihood $\mathbb{E}_{z\sim q_{\phi}(z[x)]}[-\log p_{\theta}(x|z)]$, the other CAE parts being fixed.
4.6 Calibration of VBLE predicted uncertainties
Our method yields valuable and informative uncertainties, as it will be shown in the experiments. However our variational posterior does not exactly match the true inverse problem posterior. To further improve the uncertainty prediction, we also consider the calibration strategy proposed in [54], which was shown to propose consistent error bounds for different posterior sampling methods. Specifically, this approach consists of learning, through a calibration subset, the quantiles of the true reconstruction error conditioned on an estimate of the posterior variance. Afterward, for a given pixel, the $\alpha$-quantile of the true reconstruction error given the pixel posterior variance represents an error bound of level $\alpha$. We refer to [54] for further details.
@@ 5 Experimental setting
In the experiments, we address the problem of satellite panchromatic image restoration, that is deblurring and denoising of one spectral band typically spanning wavelengths from 450nm to 850nm. In addition, we study the joint problem of restoration and super-resolution as satellite images are more and more super-resolved. For both problems, we chose the operating point of Pléiades satellites, as it constitutes a reference for very high resolution optical satellite imaging in Europe, for which we know well the degradation model. In this section, we present the considered inverse problems. Then, data collection and processing are detailed. Finally, the experiment setup is provided, including the evaluation metrics, the considered baselines, and the training details for the proposed methods.
@@ 5.1 Forward models and considered inverse problems
Deep regularization methods are model-based, thus require the log-likelihood of the forward model $p_{Y|X}(y|x)$ as it is used during the restoration process. The blurring process can hardly be exactly modelled, as the observed landscape, as well as the blur kernel, are continuous. Hence, we consider the forward model to be:
@@
where $y$ is the measure, $x_{target}$ the target image at the target resolution, $h$ the blur kernel sampled at the target resolution, $\mathcal{D}$ corresponds to a downsampling operator, and $w$ is the approximated Poisson-Gaussian noise of variance $\sigma^{2}_{w}(x_{target})=\sigma_{0}^{2}+K(\mathcal{D}(h*x_{target}))$. In the experiments, we consider two inverse problems. The first one, called IR, consists of joint denoising and deblurring. In this case, $\mathcal{D}=I$. The second one, called IR+SISR, consists of joint denoising, deblurring, and super-resolution by a factor of 2.
5.2 Detailed data processing

### v2

@@ 3 Variational Bayes joint latent-image estimation with accurate uncertainties
Figure 1: Principle of the proposed method, VBLE-xz. First, a CAE is trained on target images. Then, given a degraded image, the joint latent and image posterior is approximated following algorithm 1. Finally, posterior sampling is performed using the explicit approximate posterior, enabling to derive the posterior mean and uncertainty quantification maps.
@@ 3.1.1 Variational autoencoders
[…]
@@
3.1.2 Variational compressive autoencoders
Compressive autoencoders (CAEs) are neural networks used for compression, which have similarities to VAEs. CAEs are trained by optimizing the so-called rate-distortion trade-off.
= $\mathcal{L}=\mbox{ Rate}+\alpha\times\mbox{Distortion}.$ (7)
The distortion represents an error between the input and output of the CAE, and the rate measures the compression efficiency. CAEs can be expressed as VAEs with specific inference and generative distributions [Balle2017]. In particular, the encoder posterior $q_{\phi}(z|x)$ is a uniform distribution, of mean $\bar{z}$ and with a fixed variance in order to simulate the quantization process. Compared to VAE loss, the distortion resembles the data fidelity term, while the rate corresponds to the latent constraint. The trade-off parameter $\alpha$ plays a similar role as $\gamma$ for VAEs, controlling the trade-off between the two loss terms.
In addition, state-of-the-art CAEs use a hyperprior module [Balle2017, Balle2018, Cheng2020], as described in the first part of Fig. 1, which is learned jointly with the main autoencoder. This additional autoencoder takes $\bar{z}$ as input and predicts $z$’s mean $\mu^{z}$ and deviation $\sigma^{z}$. The latent prior on $z$ is then defined as the factorized distribution $\prod_{k}\mathcal{N}(\mu^{z}_{k},(\sigma^{z}_{k})^{2})$. Hence, the latent prior is a $z$-adaptive distribution, more powerful and flexible than the traditional VAE prior $\mathcal{N}(0,I)$. Formally, these CAEs can be expressed as VAEs with two latent variables $z$ and $h$ [Balle2018].
3.1.3 Multi-rate CAEs
To modulate the CAE compression rate, different CAEs should be trained at different values of $\alpha$ in eq. 7. However, several methods exist to design CAEs at multiple rates. The majority of these approaches [Yang2020, Song2021, Arezki2024a, Mijares2023] propose to adapt the CAE structure and then to train the resulting network with a range of $\alpha$ values. Interestingly, [Arezki2024] proposes a training-free adaptation of a CAE that was initially trained at a high bitrate. Indeed, given such a CAE, [Arezki2024] introduces a scaling parameter $s\in]0,1]$ at inference. Then, the auto-encoding process consists of:
$z\sim q_{\phi}(z|s\times x),\hat{x}=D_{\theta}(z)/s.$ (8)
@@ 3.2 Background on VBLE with CAEs
Variational Bayes Latent Estimation (VBLE) [Biquard2023] is an algorithm used for image restoration, which enables efficient approximate posterior sampling. It belongs to latent optimization methods [Bora2017], which estimate the solution of an inverse problem in the latent space of a generative model. In place of a classical generative model, [Biquard2023] proposed to use CAEs trained on image compression tasks.
Furthermore, VBLE estimates the latent inverse problem posterior $p_{Z|Y}(z|y)$ in the latent space of the CAE using variational inference. Specifically, it considers the following factorized parametric family:
$E_{\bar{z},a}$ $=\left\{q_{\bar{z},a}(z)\big|\bar{z},a\in\mathbb{R}^{C\times M\times N},a>0\right\}$ (9)
$\text{ with }q_{\bar{z},a}(z)$ $=\prod_{k}\mathcal{U}(z_{k};[\bar{z}_{k}-\frac{a_{k}}{2},\bar{z}_{k}+\frac{a_{k}}{2}])$
@@
$\mathcal{L}(\bar{z},a;y)=\mathbb{E}_{q_{\bar{z},a}(z)}\big[\log p_{Y|Z}(y|z)+\lambda(\log p_{\theta}(z)$
$-\log q_{\bar{z},a}(z))\big]$ (10)
where $\lambda$ is a hyperparameter close to 1 which is added to adapt the regularization strength, $p_{\theta}(z)$ is the latent prior, that is the hyperprior $\mathcal{N}(\mu^{z},\sigma^{z})$ in the case of CAEs, with the notation of Fig. 1. Using the reparameterization trick $z=\bar{z}+au$ with $u\sim\mathcal{U}(-\frac{1}{2},\frac{1}{2})$ [DiederikPKingma2014], eq. 10 can be optimized with a Stochastic Gradient Variational Bayes (SGVB) algorithm [DiederikP.Kingma2018].
@@ 3.3 Limitation of latent posterior estimation
Figure 2: Visual result of VBLE and VBLE-xz on an inpainting problem. MMSE is the posterior mean computed on 100 posterior samples. The error quantile is a perpixel predicted error quantile computed with 100 posterior samples. ©CNES 2025
In practice, the $\lambda$ parameter is not enough to adapt the regularization for inverse problems with very different levels of ill-posedness. A much more efficient way to control the regularization strength is the CAE bitrate, as a strong compression leads to a significant regularization. Hence, a first limitation of VBLE is the requirement to train CAEs at several bitrates to tackle inverse problems of various difficulties.
Furthermore, estimating the inverse problem uncertainty only in the latent space of the CAE works well in general but suffers from two main limitations. First, this prevents from modeling the representation error, that occurs when the solution does not lie in the decoder range. Second, the latent space posterior fails at capturing high-frequency patterns. This specific point is illustrated in Fig. 2: for an inpainting problem with a random mask, VBLE does not recover the mask in its predicted error quantile. In contrast, VBLE-xz, which we introduce in the next section, enables to identify the mask, assigning greater uncertainty to pixels that have not been observed.
3.4 VBLE-xz: Proposed joint latent and image posterior model
3.4.1 VBLE-xz algorithm
In the following, we propose a new method for image restoration, derived from VBLE, in which we model the inverse problem posterior jointly in the latent and in the image spaces.
We suppose that a CAE was trained on target images. We also assume that the generative model of this CAE is $p_{\theta}(x|z)=\mathcal{N}(D_{\theta}(z),\Sigma_{\theta}(z):=\operatorname{diag}(\sigma_{\theta}(z)^{2})$, as illustrated in Fig. 1. The training of the decoder that produces $\sigma_{\theta}(z)$ will be detailed later. For simplicity, we only consider one latent variable $z$, but the equation can easily be adapted for two latent variable CAEs by considering a concatenated variable $\psi=(z,h)$ and $p(\psi)=p(z|h)p(h)$.
Figure 3: Generative model graph of our variational inference algorithm. The latent sampling from $q_{\bar{z},a}$ and the image sampling from $q_{b}(x|z)=\mathcal{N}(x;D_{\theta}(z)/s,b\sigma_{\theta}(z)/s)$ are illustrated in the upper part of the Figure.
We aim to estimate the joint latent and image inverse problem posterior $p_{X,Z|Y}(x,z|y)$ using variational inference. Note that, for VBLE, the targeted distribution was the latent posterior $p_{Z|Y}(z|y)$. For that purpose, we consider the following parametric family, corresponding to the representation of Fig. 3:
$E_{\bar{z},a,b}$ $=\{q_{b}(x|z)q_{\bar{z},a}(z)|\bar{z},a\in\mathbb{R}^{C\times M\times N},$
@@ 3.4.1 VBLE-xz algorithm
with $C^{\prime},M^{\prime}$ and $N^{\prime}$ being the channel and spatial dimensions of the image space. The parametric distribution is composed of a latent term $q_{\bar{z},a}(z)$ and of an image term $q_{b}(x|z)$ which was designed to take into account the representation error. As in VBLE, the latent term is chosen to be uniform as it matches the CAE encoder distribution. The choice for $q_{b}(x|z)$ is motivated by the fact that $p_{\theta}(x|z)$ represents the likelihood of generating $x$ given $z$. Hence, even if the representation error for an image $x^{*}$ is nonzero (that is $D_{\theta}(z)\neq x^{*}$), $x^{*}$ should nevertheless lie in a zone where $p_{\theta}(x|z)$ is high. Then, parameter $b$ allows to better adapt the posterior model to the given inverse problem. This idea is visually illustrated in Fig. 3.
@@
$+\lambda(\log p_{\theta}(x|z)+\log p_{\theta}(z)-\log q_{b}(x|z)-\log q_{\bar{z},a}(z))\big]$ (11)
with $p_{\theta}(z)$ being the latent prior of the generative model, that is the hyperprior in the CAE case, and $\lambda$ a hyperparameter close to one, added to control the regularization as in VBLE. Then, the $(x|z)$ terms represent a KL divergence between two Gaussian distributions and can be easily simplified, as well as $\log q_{\bar{z},a}(z)$ which is independent of $z$. Thus:
@@
$+\lambda\log p_{\theta}(z)\big]+\lambda\left(\sum_{k}\log a_{k}+\sum_{i}\log b_{i}-\frac{1}{2}b_{i}^{2}\right).$ (12)
@@
eq. 12 can be optimized using gradient descent with a Monte Carlo approximation of the ELBO gradient as in [DiederikPKingma2014, Biquard2023].
Algorithm 1 Posterior approximation with VBLE-xz
$y$ measure, ($\lambda,s$) hyperparameters, $\bar{z}_{0}\in\mathbb{R}^{C\times M\times N}$, $a_{0}\in\mathbb{R}^{C\times M\times N}=(1)_{i,j,l}$, $b_{0}\in\mathbb{R}^{C^{\prime}\times M^{\prime}\times N^{\prime}}=(1)_{i,j,l}$ $l=0$, $\eta>0$, sgsg stop-gradient operation.
while not convergence do
  $z=\bar{z}_{l}+a_{l}u$ with $u\sim\prod_{k}\mathcal{U}(-0.5,0.5)$
  $x=D_{\theta}(z)/s+(b_{l}sg(\sigma_{\theta}(z))/s)\epsilon$ with $\epsilon\sim\mathcal{N}(0,I_{C^{\prime}\times M^{\prime}\times N^{\prime}})$
  $\begin{pmatrix}\bar{z}_{l+1}\\ a_{l+1}\\ b_{l+1}\end{pmatrix}=\begin{pmatrix}\bar{z}_{l}\\ a_{l}\\ b_{l}\end{pmatrix}+\eta\nabla_{\bar{z},a,b}\big[\log p_{Y|X}(y|x)+\lambda(\log p_{\theta}(z)+\sum_{k}\log a_{l,k}+\sum_{j}\log b_{l,j}-\frac{1}{2}b_{l,j}^{2})\big]$
  $l=l+1$
end while
return $\bar{z}_{l},a_{l},b_{l}$
3.4.2 Modulating the CAE bitrate in VBLE-xz
We introduce an additional scaling parameter in VBLE-xz in order to modulate the CAE bitrate in the same manner of eq. 8 [Arezki2024]. This scaling enables to modulate the regularization of the inverse problem without having to train several CAEs at several bitrates. Let $s\in]0,1]$ be a hyperparameter. Introducing the scaling into VBLE-xz amounts to change the distributions over $(x|z)$ in the following way:
$p_{\theta}(x|z)=\prod_{k}\mathcal{N}(D_{\theta}(z)_{k}/s,\sigma_{\theta}(z)_{k}^{2}/s^{2}),$
$q_{b}(x|z)=\prod_{k}\mathcal{N}(D_{\theta}(z)_{k}/s,b_{k}^{2}\sigma_{\theta}(z)_{k}^{2}/s^{2}).$ (14)
It induces slight changes in the sampling procedure described in eq. 13. Apart from this, the loss in equations eq. 11 and eq. 12 stays unchanged. The whole VBLE-xz algorithm is summarized in algorithm 1.
3.4.3 Learning the generator variance
The algorithm 1 requires the following generative model: $p_{\theta}(x|z)=\mathcal{N}(D_{\theta}(z),\text{diag}(\sigma_{\theta}(z)^{2})$, thus the generator should estimate a pixelwise variance $\sigma_{\theta}(z)$, as illustrated in Fig. 1. However, in classical VAE and CAE models, the decoder distribution is a Gaussian distribution $p_{\theta}(x|z)=\mathcal{N}(D_{\theta}(z),\gamma^{2}I)$ with a fixed perpixel standard deviation $\gamma$ responsible for the trade-off between the data fidelity and the latent constraint terms in the loss. This distribution is not appropriate in our case as the reconstruction error should be larger on high frequency areas than on large flat ones.
Some approaches learn more representative generator distributions [Rybkin2021, Duff2023, Dorta2018]. Following [Rybkin2021], we add a second decoder predicting $\sigma_{\theta}(z)$, which has the same structure as the original decoder $D_{\theta}$. We adopt a two stage training strategy as is typically done for training stability: first, the CAE is classically trained with a given $\gamma$ trade-off parameter. Then, the second decoder is trained by minimizing the negative log-likelihood $\mathbb{E}_{z\sim q_{\phi}(z[x)]}[-\log p_{\theta}(x|z)]$, the other CAE parts being fixed.
During algorithm 1, we introduce a stop-gradient operation on the deviation map $\sigma_{\theta}(z)$, that is $\sigma_{\theta}(z)$ is treated as a constant with respect to $z$ at each iteration. Without this operation, we observed that the deviation map abnormally converged toward zero, which prevented the $b_{k}$ coefficients from being correctly estimated. This behavior is likely due to the strong over-parameterization of the latent space in CAEs. Since addressing this issue during CAE training is not straightforward, the stop-gradient constitutes a relevant alternative that fully resolves the problem.
@@ 4 Experimental setting
In the experiments, we consider the processing of one panchromatic spectral band, typically spanning wavelengths from 450nm to 850nm. We choose the operating point of Pléiades satellites with a resolution of 50cm, as it constitutes a reference for very high-resolution optical satellite imaging in Europe, for which we have a very precise model of the degradation. In this section, we present the considered inverse problems. Then, data collection and processing are detailed. Finally, the experiment setup is provided, including the evaluation metrics, the considered baselines, and the training details for the proposed methods.
@@ 4.1 Forward models and considered inverse problems
We consider three inverse problems: image restoration (IR), that is deblurring and denoising, joint image restoration and super-resolution (IR+SISR), and inpainting, with random or structured masks. The inpainting problem is noiseless. For IR and IR+SISR, the blurring process can hardly be exactly modelled, as the observed landscape, as well as the blur kernel, are continuous. Hence, we consider the forward model to be:
@@
where $y$ is the measure, $x_{target}$ the target image at the target resolution, $h$ the blur kernel sampled at the target resolution, $\mathcal{D}$ corresponds to a downsampling operator, and $w$ is the approximated Poisson-Gaussian noise of variance $\sigma^{2}_{w}(x_{target})=\sigma_{0}^{2}+K(\mathcal{D}(h*x_{target}))$. For IR, $\mathcal{D}=I$ and for IR+SISR $\mathcal{D}$ is a decimation operator by a factor of 2.
4.2 Detailed data description